# Multiscale Hodge Scattering Networks for Data Analysis

### v4

@@ 6 Domain Classification
Another vital application of geometric scattering networks and graph neural networks (GNNs) is graph (and simplex) classification. Broadly speaking, this problem consists of predicting a label of a social or chemical graph based on a training set of similar graphs with different configurations (i.e., different numbers of nodes and edges). Since the size of the graphs varies greatly within these datasets, we employ only the global-pooling version of our MHSN, akin to the previous efforts reported in [4, 25], which were based on geometric scattering methods. We employ a simple SVM with Gaussian radial basis functions to classify the features extracted by the MHSTs. Moreover, we tune the hyperparameters controlling the strength of the $\ell^{2}$-regularization and the kernel coefficients via the cross-validation scheme presented in [4] using the same search space and validation indexes.
In some applications, there is additional labeling (node labels, edge labels, etc.), but we consider the unlabeled case for now. Therefore, we have to compute permutation-invariant input features based only on topological/geometric information from the boundary matrices. Since many of the graphs are too small to contain high-degree simplices, we only consider node and edge-based features and transforms. For the node-based signals, we first compute the eccentricity and clustering coefficient of each node. Then, pass each into the scattering transform and concatenate the outputs (using global pooling). Finally, we used these features as inputs to an SVM. For the edge-based signals, we use two sets of features: 1) the number of off-diagonal terms in Hodge Laplacian; and 2) the average degree of the head and tail nodes of each edge. For each problem we train three models: 1) using only node features; 2) using only edge features; and 3) using both sets. We adopt Gaussian radial basis functions with hyperparameter space and tuning scheme proposed by [4].
@@ 7 Molecular Dynamics
We extract a rich set of features for each structure (i.e., a pose or conformation of a molecule) using our MHSNs and then employ a support vector regression (SVR) method with Gaussian radial basis functions to approximate and predict the energy. More specifically, for each molecule, we first compute the average position of each atom across the training set. Then, using these positions, we create a $k$NN-graph (with $k=5$) as a template simplicial complex. Note that by using this simplicial complex, rather than the molecular-bond graph, we can better capture the geometric information in the pose of the molecule as detailed in [47, 48]. We then generate the node and edge signals of both training and test sets as follows. We compute the Euclidean distance matrix of the node coordinates of each snapshot. For each node in the simplicial complex, we assign the corresponding column vector of the distance matrix as its node value. This generates a vector-valued node signal per snapshot. As for an edge signal for a snapshot, we extract edge lengths from the above distance matrix and create a diagonal matrix of edge lengths. Then, we assign the corresponding column vector of the edge distance matrix as its edge value. This generates a vector-valued edge signal per snapshot. We do not use simplices of dimension greater than $1$ because the molecules are too small.

### v5

@@ 6 Domain Classification
Another vital application of geometric scattering networks and graph neural networks (GNNs) is graph (and simplex) classification. Broadly speaking, this problem consists of predicting a label of a social or chemical graph based on a training set of similar graphs with different configurations (i.e., different numbers of nodes and edges). For example, in the COLLAB dataset [52], each graph represents a network of coauthors of a scientific paper. Since the size of the graphs varies greatly within these datasets, we employ only the global-pooling version of our MHSN, akin to the previous efforts reported in [4, 25], which were based on geometric scattering methods.
We compute permutation-invariant input features based only on topological information obtainable from the boundary matrices. Since many of the graphs are too small to contain high-degree simplices, we only consider node and edge-based features and transforms. Following the methodology developed in [4], we set $(J,M,Q)=(4,2,4)$. For the node signals, we first compute the eccentricity and clustering coefficient [HARRIS-ETAL, Sec. 1.2] of each node. For each node signal, the number of parameters (MHST coefficients) are 64 via the formula $Q\sum_{k=0}^{M}\binom{J+1}{k}$, hence 128 parameters after concatenating them. For the edge signals, we use the average of the eccentricities of the head and tail nodes of each edge and the number of non-zero off-diagonal terms in the combinatorial Hodge-Laplacian (each such term corresponds to a $1$-adjacent edge [29, Sec. 4.1]).
For each domain classification problem we train three models: 1) using 128 node features; 2) using 128 edge features; and 3) using 256 combined features. We then employ a simple SVM with Gaussian radial basis functions to classify the features. Moreover, we tune the hyperparameters controlling the strength of the $\ell^{2}$-regularization and the kernel coefficients via the cross-validation scheme presented in [4] using the same search space and validation indexes.
@@ 7 Molecular Dynamics
We extract a rich set of features for each structure (i.e., a pose or conformation of a molecule) using our MHSNs (without pooling) and then employ a support vector regression (SVR) method with Gaussian radial basis functions to approximate and predict the energy. More specifically, for each molecule, we first compute the average position of each atom across the training set. Then, using these positions, we create a $k$NN-graph (with $k=5$) as a template simplicial complex. Note that by using this simplicial complex, rather than the molecular-bond graph, we can better capture the geometric information in the pose of the molecule as detailed in [47, 48]. Unlike the domain classification problems in Section 6, the geometry of the simplicial complex is fixed, so rather than use its geometrically-invariant descriptors, we need to begin with signals that encode the position information of molecules. We then generate the node and edge signals of both training and test sets as follows. First we compute the Euclidean distance matrix of (i.e the Gram matrix of the point-cloud, measured in the Euclidean distance) the node coordinates of each snapshot and assign the corresponding column vector of the distance matrix as its node features. This generates a number-of-atoms-dimensional node signal for each node, for each snapshot. For an edge signal, we extract edge lengths from the above distance matrix and create a diagonal matrix of edge lengths. Then, we assign the corresponding column vector of this diagonal matrix as its edge features. As with the node-based signal, this gives us number-of-edges-dimensional signal to input into our MHSN. For this experiment we use $(J,M,Q)=(4,2,3)$. We do not use simplices of dimension $\kappa>1$ and not set $J>4$ and $Q>3$ because the molecules are too small.